Meiston Press has a debt-equity ratio of 1.80. The pre-tax cost of debt is 9.00 percent and the cost of equity is 14.1 percent. What is the firm’s weighted average cost of capital (WACC) if the tax rate is 34 percent?

  • 10.00 percent

  • 8.85 percent

  • 9.63 percent

  • 10.88 percent
Answer #1

After-tax cost of debt=9*(1-tax rate)

=9(1-0.34)=5.94%

Debt-equity ratio=debt/equity

Hence debt=1.8*equity

Let equity be $x

Debt=$1.8x

Total=$2.8x

WACC=Respective costs*Respective weight

=(x/2.8x*14.1)+(1.8x/2.8x*5.94)

=8.85%(Approx).